Grade 12th PassMechanics

a uniform elastic string has length a1 when tension is t1 and length a2 when tension is t2 . the amount of work done in streching it from its natural length to length a1+a2

sana samad
14 Years agoGrade 12th Pass

1 Answer

Chetan Mandayam Nayakar
14 Years ago

Let a0 be natural length and k be stiffness of string

a0 +(t1/k)=a1

a0 +(t2/k)=a2

solve for a0 and k

answer=(k/2)(a1+a2-a0)2